# A practical identifiability criterion leveraging weak-form parameter estimation

### v2

@@ 4.2 Results of practical identifiability
We estimate the parameters $w_{1},$ $w_{2}$, and $w_{3}$ from the weak-form input-output equation (5) using WENDy (see Appendix A) for simulated data with Gaussian additive error. We vary the measurement additive error ratio in the simulated data between $e\in[0\%,20\%]$ such that the standard deviation of the additive error is $\sigma=e\text{RMS}(\Omega(t))$, and at each additive error ratio, we generate 1,000 noisy data sets. The (e,q)-identifiability (see Section 2.2) for $e\in[0\%,20\%]$ and $q\in[1\%,100\%]$ of model (5) is given in Figure 4A, where the area in blue represents where the model is (e,q)-identifiable. For example, the model is (5,50)-identifiable, meaning that at a $5\%$ additive error ratio in the data, the MSE of the parameter estimates $\widehat{w}=[\widehat{w}_{1},\widehat{w}_{2},\widehat{w}_{3}]$ falls below $50\%\times w_{i}$. However, at $15\%$ measurement additive error ratio, this property no longer holds, i.e., the model is not (15,50)-identifiabile. We also observe from the $(e,q)-$identifiability of the individual parameters that the identifiability parameter $w_{3}$ is lost first as noise increases. In the example above, both $w_{2}$ and $w_{3}$ are $(15,50)-$identifiable, while $w_{3}$ is not. As demonstrated in Figure 4A, as additive error increases, the MSE increases and approaches the magnitude of the true parameters.
@@
Figure 4: Using WENDy with equation (5) we find the blood concentration model to be generally practically identifiable below $11\%$ scaled noise in the data. (A.) The area in blue denotes where the model is (e,q)-identifiable, and the area in white denotes where the model is not (e,q)-identifiable. The (e,q)-identifiability was determined from 1,000 samples at each respective error level, e. The black $\boldsymbol{\mathsf{X}}$ denotes the $(5,50)-$identifiability cutoff. Because this point falls within the blue region, we can say that at a $5\%$ additive error ratio in the data the MSE of $\widehat{w}_{1},$ $\widehat{w}_{2}$, and $\widehat{w}_{3}$ remain below the square of $50\%$ the parameter magnitude. (B.) The relative error determined from 1,000 samples at each respective error level, e. (C.) The proportion of $95\%$ confidence intervals at $e=[0\%,20\%]$ that contain the true value of $w_{1},$ $w_{2},$ and $w_{3}$.
@@
We additionally compare the (e,q)-identifiability of model (5) to the average relative error and the coverage of the $95\%$ confidence intervals generated at each additive error ratio. Previous work by our group and others [44, 19] proposed (over many data sets generated at a particular additive error ratio) that a model can be considered practically identifiable in terms of relative error, if the relative error is equal to or lower than the error introduced through additive noise. The relative error generated for 1,000 simulated data sets at each additive error level is given in Figure 4B, and the coverage of the $95\%$ confidence intervals for $\widehat{w}_{1}$, $\widehat{w}_{2}$, and $\widehat{w}_{3}$ are given in Figure 4C. Based on this strict relative error criterion, the blood concentration model is not practically identifiable for any error level. Even at $1\%$ additive noise, the relative error in parameter $w_{3}$ is $1.1\%$. However, a more relaxed condition on relative error for practical identifiability has been used in [19], where relative error needs only to be below $10e$. In this case, the blood concentration model is practically identifiable with respect to the relative error from $e=[0\%,20\%]$. Generally, the relative error remains near $2\times e$ for parameters $w_{1}$ and $w_{2}$ and near $3\times e$ for parameter $w_{3}$ for $e=[0\%,20\%]$. Additionally, above the $11\%$ additive error level, where the model is no longer (e,50)-identifiable, the coverage for all parameters $w_{1}$, $w_{2}$, and $w_{3}$ decreases as shown in Figure 4C. For parameters $w_{1}$ and $w_{2}$ at error levels below $e=6\%$ in all trials, the $95\%$ confidence interval contains the true values of $w_{1}$ and $w_{2}$. This decreases below $70\%$ at $e=11\%$ measurement additive error ratio. All three criteria agree that parameter $w_{3}$ becomes non-identifiable first as noise increases.
@@
We estimate the parameter $\beta$ from the modified weak-form input-output equations (13) using WENDy for simulated data with Gaussian additive error. We vary the additive error ratio in the simulated data between $e=[0\%,200\%]$ such that the standard deviation of the additive error is $\sigma=e\text{RMS}(\Omega(t))$, and at each additive error ratio, we generate 1,000 noisy data sets. The (e,q)-identifiability for $e=[0\%,200\%]$ and $q=[1\%,100\%]$ of model (13) is given in Figure 5A. We find that the weak form SIR model is (e,q)-identifiable for all estimate error ratios above the $q=10\%$ level. For measurement error levels from $[0\%,120\%]$ the MSE always remains below $5\%\times\beta$. Therefore, we conclude the $\beta$ is generally practically identifiable for noise up to $120\%\times\text{RMS}(\Omega(t))$.
@@
Figure 5: Using WENDy with equations (13) we find the SIR model to be generally practically identifiable for additive noise between $[0\%,200\%]$. (A.) The area in blue denotes where the model is (e,q)-identifiable, and the area in white denotes where the model is not (e,q)-identifiable. The (e,q)-identifiability was determined from 1,000 samples at each respective error level, e. The black $\boldsymbol{\mathsf{X}}$ denotes the $(50,10)-$identifiability cutoff. Because this point falls within the blue region, we can say that at a $50\%$ additive error ratio in the data, the MSE of $\widehat{\beta}$ remains below the square of $10\%$ of the magnitude of $\beta$. (B.) The relative error determined from 1,000 samples at each respective error level, e. (C.) The proportion of estimated $95\%$ confidence intervals at $e=[0\%,200\%]\%$ that contain the true value of $\beta$.
@@
We additionally compare the (e,q)-identifiability of model equations (11)-(12) to the average relative error and the coverage of the $95\%$ confidence intervals generated at each additive error ratio. The relative error generated for 1,000 simulated data sets at each additive error level is given in Figure 5B, and the coverage of the $95\%$ confidence intervals for $\beta$ is given in Figure 5C. Here, the (e,q)-identifiability and the relative error criteria are in agreement. The relative error remains below $6\%$ even at high additive error ratio, $e=200\%$. Unlike the blood concentration model, the coverage remains $100\%$ across all additive error ratios. However, it is important to note that while the parameter estimates remain very accurate at high noise, the uncertainty in the estimate still increases with increasing additive noise ($e$) similar to the blood concentration model.
@@
Blood-tissue diffusion $19$ sec $70$ sec
@@
Table 1: Median time to generate $N=1,000$ samples using WENDy method vs OE methods at a $5\%$ additive noise ratio for the Blood-tissue diffusion and SIR models.
Figure 6: Using WENDy with equations (5) we are able to recover the parameters $w=[w_{1},w_{2},w_{3}]$ for the blood-tissue diffusion model from noisy data both more quickly and more consistently than using an output error method. The relative error in $w$ versus walltime is given for 50 fittings of equations (5) to $x_{1}$ compartment data of varying additive error ratios found using WENDy (blue dots) or 50 fittings of model equations (4) using output error (purple diamonds) for additive error ratio between $e=[0\%,10\%]$. Note that for $\sim 60\%$ of initial guesses, the OE method fails to converge, and that only successful optimizations for the OE method have been presented above.
In Figure 7, the relative errors in parameter $\beta$ for 50 simulated data sets of additive error ratios $e=0\%,5\%,10\%,15\%,$ and $20\%$ are plotted against the walltime using WENDy on equations (13) and using an output error method on (9). At all error levels, WENDy is multiple orders of magnitude faster than the output error method. The median walltimes for the WENDy method fall between 6.4-9.1×$\times 10^{-4}$ seconds, while the median output error walltimes fall between 1.3-1.6×$\times 10^{-1}$ seconds. Table 1 shows that the WENDy method can be used in under a second to test practical identifiability on a single error ratio level. The median relative error is similar, within $2\times 10^{-3}$, between the WENDy method and the OE method at error levels $e=5\%,10\%,15\%$. However, at the $e=20\%$ additive error ratio, the relative error using WENDy is lower, $4.6\times 10^{-3},$ compared to the relative error using the output error, $1.6\times 10^{-2}.$ Ultimately, the benefit of smaller walltimes for estimation of parameters in both the blood-tissue diffusion and SIR models allows for the calculations on thousands of simulated data sets, like those used to generate Figures 4 and 5.
Figure 7: Using WENDy with equations (13) we are able to recover the transmission rate $\beta$ for the SIR model from noisy data, both more quickly than using an output error method. The relative error in $\beta$ versus walltime is given for 50 fittings of equations (13) to infected compartment data of varying additive error ratio found using WENDy (blue dots) or 50 fittings of model equations (9) using output error (purple diamonds) for additive error ratio between $e=[0\%,20\%$].
@@ 5 Conclusion
    We apply differential elimination methods to generate an input-output equation. At this step, we can also assess the structural identifiability of the model, which is prerequisite to practical identifiability.
@@
    We cast the input-output equation into its weak-form. This step involves identifying terms of the weak-form input-output equation where integration by parts can and cannot be applied. Terms where integration by parts is not possible may be managed by either a substitution of variables or an approximation of higher order derivatives.
@@
    We employ a sampling-based method of assessing practical identifiability. In this process, we generate simulated datasets for various levels of measurement error ratios, $e$. We then apply WENDy to weak-form input-ouput equations for efficient parameter estimation.
@@
The $(e,q)$-identifiability criterion offers a method for practical identifiability which based on the noise in observed variables $e$ and the mean-square error of the parameter estimator $q$. We demonstrate this $(e,q)$ criterion via computing the practical identifiability of these two examples in terms of mean-squared error and in terms of average relative error and coverage. This criterion is better able to capture changes in the quality of the parameter estimator due to increased noise in the data compared to criteria based on average relative errors. Additionally, while in this study we apply the criterion to ODE systems with Gaussian measurement error, it can be extended to a broad range of problems, including discrete-time systems, PDE systems, and problems with non-Gaussian noise, or even multiplicative rather than additive noise. [^4] Recently, WENDy has been shown to successfully recover parameters from data with multiplicative log-normal noise [45]. Assessing identifiability in PDE systems is a challenging problem, which has been addressed recently by [46, 47, 48], but it still remains difficult (if not impossible) to test structural identifiability for arbitrary PDE systems. Therefore, methods that can efficiently assess the practical identifiability of these systems are particularly important.
@@
Critically, the practical identifiability analyses in this study rely on the efficiency of weak-form parameter estimation. We present a method for weak-form parameter estimation of ODE systems with unobserved compartments, utilizing differential elimination to generate weak-form input-output equations and subsequently applying WENDy. This method is efficient, accurate, and robust to noise. Additionally, applied to a blood-tissue diffusion system (Model (4)), a comparison with the results in [1] yields that the WENDy is better than the Hartley and polynomial test functions and strongly competitive with the interpolated polynomials as well as the pure integral approach.
@@
Regarding method performance, we note that in all the examples presented above (including at all error levels), WENDy is substantially faster than OE methods. In particular, WENDy is faster than the output error method for both the blood-tissue diffusion and SIR models, and is multiple orders of magnitude faster in the SIR case. Ultimately, the benefit of smaller walltimes for estimation of parameters in both the blood-tissue diffusion and SIR models allows for the calculations on thousands of simulated data sets, like those used to generate Figures 4 and 5. Table 1 demonstrates that for a single error ratio level the WENDy method is able to generate a sample of size $N=1,000$ faster than OE methods, with a particularly large advantage (seconds vs minutes) in the SIR model case. For the SIR model the accuracy of the weak-form method is the same as the OE method for additive noise ratios between $[0\%,15\%]$, and more accurate at high noise ratios of $e=20\%$. We further observe that the relative error increases linearly with the additive error ratio for this system. Linear error has also been observed for some systems in [49, 25, 50] when performing EE regression-based optimization. While the median accuracy of the weak-form method is less than the output error method for the blood-tissue diffusion model, the weak-form method does not fail to converge (unlike OE methods, which rely on inspired initial parameter estimates). Notably, for Example 3.1, $\sim 60\%$ of the OE optimizations failed to converge, even with reasonable initial parameter guesses (see Figure 6).
@@
One limitation of this study is that in the generation of the weak-form input-output equation, information about the initial conditions of the system is lost. There are examples of ODE systems where initial conditions can play an important role in the identifiability of key model parameters [34, 51, 52]. A possible direction to address this limitation is to work with an integral representation of the equation near the end points of the domain to recover information about the initial condition lost in the transformation to the weak form. Future exploration of the impact of initial conditions and methods to incorporate this lost information in our weak-form framework is needed.
@@
Lastly, in this study, we approach practical identifiability through simulated data, and our $(e,q)-$criterion provides an a priori method of studying model identifiability and robustness to noise. However, it is also of interest to find methods of assessing the identifiability given a specific experimental data set. A future direction of this work could be to explore the relationship between experimental data and the stability of identifiability calculations.

### v3

@@ 4.2 Results of practical identifiability
We estimate the parameters $w_{1},$ $w_{2}$, and $w_{3}$ from the weak-form input-output equation (9) for the blood diffusion model using WENDy (see Section 3.2), for simulated data with Gaussian additive observation error. We vary the additive observation error ratio in the simulated data between $e\in[0\%,20\%]$ such that the standard deviation of the additive error is $\sigma=e\text{RMS}(\Omega(t))$, and at each additive observation error ratio, we generate 1,000 noisy datasets. The (e,q)-identifiability (see Section 2.2) for $e\in[0\%,20\%]$ and $q\in[1\%,100\%]$ of model (9) is given in Figure 4A, where the area in blue represents where the model is (e,q)-identifiable. For example, the model is (5,50)-identifiable, meaning that at a $5\%$ additive observation error ratio in the data, the MSE of the parameter estimator $\widehat{w}=[\widehat{w}_{1},\widehat{w}_{2},\widehat{w}_{3}]$ falls below $50\%\times w_{i}$. However, at $15\%$ additive observation error ratio, this property no longer holds, i.e., the model is not (15,50)-identifiable. We also observe from the $(e,q)$-identifiability of the individual parameters that the identifiability parameter $w_{3}$ is lost first as noise increases. In the example above, both $w_{2}$ and $w_{3}$ are $(15,50)$-identifiable, while $w_{3}$ is not. As demonstrated in Figure 4A, as additive observation error increases, the MSE increases and approaches the magnitude of the true parameters. We do not find the blood-tissue diffusion model to be generally practically identifiable, using a (10,20)-identifiable criterion.
@@
Figure 4: Using WENDy with equation (9) we find the blood concentration model to be generally practically identifiable below $11\%$ scaled noise in the data. (A.) The area in blue denotes where the model is (e,q)-identifiable, and the area in white denotes where the model is not (e,q)-identifiable. The (e,q)-identifiability was determined from 1,000 samples at each respective error level, e. The black $\boldsymbol{\mathsf{X}}$ denotes the $(5,50)$-identifiability cutoff. Because this point falls within the blue region, we can say that at a $5\%$ additive observation error ratio in the data the MSE of $\widehat{w}_{1},$ $\widehat{w}_{2}$, and $\widehat{w}_{3}$ remain below the square of $50\%$ the parameter magnitude. The red star marks the $(10,20)$-identifiability criterion. (B.) The relative error determined from 1,000 samples at each respective error level, e. (C.) The proportion of $95\%$ confidence intervals at $e\in[0\%,20\%]$ that contain the true value of $w_{1},$ $w_{2},$ and $w_{3}$.
@@
We additionally compare the (e,q)-identifiability of model (9) to the average relative error and the coverage of the $95\%$ confidence intervals generated at each additive observation error ratio. Previous work by our group and others [48, 21] proposed (over many datasets, $D\geq 1,000$, generated at a particular additive observation error ratio) that a model can be considered practically identifiable in terms of relative error, if the relative error is equal to or lower than the error introduced through additive noise. The relative error generated for 1,000 simulated datasets at each additive observation error level is given in Figure 4B, and the coverage of the $95\%$ confidence intervals for ${w}_{1}$, ${w}_{2}$, and ${w}_{3}$ are given in Figure 4C. Based on this strict relative error criterion, the blood concentration model is not practically identifiable for any error level. Even at $1\%$ additive noise, the relative error in parameter $w_{3}$ is $1.1\%$. However, a more relaxed condition on relative error for practical identifiability has been used in [21, 38], where relative error needs only to be below $10e$. In this case, the blood concentration model is practically identifiable with respect to the relative error from $e=[0\%,20\%]$. Generally, the relative error remains near $2\times e$ for parameters $w_{1}$ and $w_{2}$ and near $3\times e$ for parameter $w_{3}$ for $e=[0\%,20\%]$. Additionally, above the $11\%$ additive observation error level, where the model is no longer (e,50)-identifiable, the coverage for all parameters $w_{1}$, $w_{2}$, and $w_{3}$ decreases as shown in Figure 4C. For parameters $w_{1}$ and $w_{2}$ at error levels below $e=6\%$ in all trials, the $95\%$ confidence interval contains the true values of $w_{1}$ and $w_{2}$. This decreases below $70\%$ at $e=11\%$ additive observation error ratio. All three criteria agree that parameter $w_{3}$ becomes non-identifiable first as noise increases.
Example 4.2.
@@
We estimate the parameter $\beta$ from the modified weak-form input-output SIR equations (17) using WENDy for simulated data with Gaussian additive error. We vary the additive error ratio in the simulated data between $e=[0\%,200\%]$ such that the standard deviation of the additive error is $\sigma=e\text{RMS}(\Omega(t))$, and at each additive error ratio, we generate 1,000 noisy datasets. The (e,q)-identifiability for $e=[0\%,200\%]$ and $q=[1\%,100\%]$ of model (17) is given in Figure 5A. We find that the weak form SIR model is (e,q)-identifiable for all estimate error ratios above the $q=10\%$ level. For measurement error levels from $0\%,120\%]$ the MSE always remains below $5\%\times\beta$. Therefore, we conclude the $\beta$ is generally practically identifiable for noise up to $120\%\times\text{RMS}(\Omega(t))$.
@@
Figure 5: Using WENDy with equations (17) we find the SIR model to be generally practically identifiable for additive noise between $0\%,200\%]$. (A.) The area in blue denotes where the model is (e,q)-identifiable, and the area in white denotes where the model is not (e,q)-identifiable. The (e,q)-identifiability was determined from 1,000 samples at each respective error level, e. The black $\boldsymbol{\mathsf{X}}$ denotes the $(50,10)$-identifiability cutoff. Because this point falls within the blue region, we can say that at a $50\%$ additive error ratio in the data, the MSE of $\widehat{\beta}$ remains below the square of $10\%$ of the magnitude of $\beta$. The red star marks the $(10,20)$-identifiability criterion. (B.) The relative error determined from 1,000 samples at each respective error level, e. (C.) The proportion of estimated $95\%$ confidence intervals at $e=[0\%,200\%]\%$ that contain the true value of $\beta$.
@@
We additionally compare the (e,q)-identifiability of model equations (17) to the average relative error and the coverage of the $95\%$ confidence intervals generated at each additive error ratio. The relative error generated for 1,000 simulated datasets at each additive error level is given in Figure 5B, and the coverage of the $95\%$ confidence intervals for $\beta$ is given in Figure 5C. Here, the (e,q)-identifiability and the relative error criteria are in agreement. The relative error remains below $6\%$ even at a high additive error ratio, $e=200\%$. Unlike the blood concentration model, the coverage remains $100\%$ across all additive error ratios. However, it is important to note that while the parameter estimates remain very accurate at high noise, the uncertainty in the estimate still increases with increasing additive noise ($e$), similar to the blood concentration model.
Additionally, we consider an example of (e,q)-identifiability where noise is multiplicatively lognormally distributed, i.e., $y(t)=\Omega(U(t),w)\epsilon$ and $\epsilon\sim\text{logNorm}(0,e).$ We estimate the parameter $\beta$ using WENDy for simulated data with log-normal multiplicative observation error. We vary the multiplicative observation error ratio in the simulated data between $e=0\%$ and $e=20\%$ such that the standard deviation of the error is $\sigma=e\log(\text{RMS}(\Omega(t)))$, and at each multiplicative observation error ratio, we generate 1,000 noisy datasets. The (e,q)-identifiability for $e=[0\%,20\%]$ and $q=[1\%,100\%]$ of model (17) is given in Figure 6A. Unlike in the case of additive noise, we find that the weak form SIR model is (e,q)-identifiable for all estimate error ratios above the $q=40\%$ level, but not the $q=10\%$ level. At error ratios above $e=16\%$, the SIR model is no longer (e,10)-identifiable. Multiplicative noise causes a greater impact near the peak in the SIR model, which we suspect accounts for the difference in parameter recovery between the additive and multiplicative noise cases. Note that the model is still $(10,20)$-identifiable. Therefore, we consider a simple SIR model with daily observations to be generally practically identifiable for both additive Gaussian and multiplicative lognormal noise.
Figure 6: Using WENDy with equations (17) we find the SIR model to be generally practically identifiable for multiplicative noise in $0\%,20\%]$. (A.) The area in blue denotes where the model is (e,q)-identifiable, and the area in white denotes where the model is not (e,q)-identifiable. The (e,q)-identifiability was determined from 1,000 samples at each respective error level, e. The red star marks the $(10,20)$-identifiability criterion. (B.) The relative error determined from 1,000 samples at each respective error level, e. (C.) The proportion of estimated $95\%$ confidence intervals across $e=[0\%,20\%]$ that contain the true value of $\beta$.
We additionally compare the (e,q)-identifiability of model equations (17) to the average relative error and the coverage of the $95\%$ confidence intervals generated at each multiplicative error ratio. The relative error generated for 1,000 simulated datasets at each multiplicative observation error level is given in Figure 6B, and the coverage of the $95\%$ confidence intervals for $\beta$ is given in Figure 6C. Here, the (e,q)-identifiability and the relative error criteria are in agreement. The relative error remains below $12\%$ for multiplicative observation error ratios $e=[0\%,20\%]$. Similarly to the case with Gaussian additive noise, the coverage remains $100\%$ across all additive error ratios. Again, it is important to note that while the parameter estimates remain accurate at high noise, the uncertainty in the estimate still increases with increasing additive noise ($e$).
@@
Blood-tissue diffusion 1919 sec $70$ sec
@@
Table 1: Median time to generate parameter estimates for $D=1,000$ simulated datasets using the WENDy method vs the OE methods at a $5\%$ additive noise ratio for the Blood-tissue diffusion and SIR models.
Figure 7: Using WENDy with equations (9) we are able to recover the parameters $w=[w_{1},w_{2},w_{3}]$ for the blood-tissue diffusion model from noisy data both more quickly and more consistently than using an output error method. The relative error in $w$ versus walltime is given for 50 fittings of equations (9) to $x_{1}$ compartment data of varying additive error ratios found using WENDy (blue dots) or 50 fittings of model equations (7) using output error (purple diamonds) for additive error ratio in $e=[0\%,10\%]$. Note that for $\sim 60\%$ of initial guesses, the OE method fails to converge, and that only successful optimizations for the OE method have been presented above. In contrast, WENDY converges in all datasets.
[…]
Figure 8: Using WENDy with equations (17), we are able to recover the transmission rate $\beta$ for the SIR model from noisy data, both more quickly than using an output error method. The relative error in $\beta$ versus walltime is given for 50 fittings of equations (17) to infected compartment data of varying additive error ratio found using WENDy (blue dots) or 50 fittings of model equations (13) using output error (purple diamonds) for additive error ratio in $e=[0\%,20\%$].
@@ 5 Conclusion
    We apply differential elimination methods to generate an input-output equation. At this step, we can also assess the structural identifiability of the model, which is a prerequisite to practical identifiability.
@@
    We cast the input-output equation into its weak form. This step involves identifying terms of the weak-form input-output equation where integration by parts can and cannot be applied. Terms where integration by parts is not possible may be managed by either a substitution of variables or an approximation of higher-order derivatives.
@@
    We employ a simulation-based method of assessing practical identifiability. In this process, we generate simulated datasets for various levels of observational error ratios, $e$. We then apply WENDy to weak-form input-output equations for efficient parameter estimation.
@@
The $(e,q)$-identifiability criterion offers a method for assessing practical identifiability which is based on the noise in observed variables $e$ and the mean-square error of the parameter estimator $q$. We demonstrate this $(e,q)$ criterion via computing the practical identifiability of these two examples in terms of mean-squared error and in terms of average relative error and coverage. This criterion is better able to capture changes in the quality of the parameter estimator due to increased noise in the data compared to criteria based on average relative errors. Additionally, while in this study we apply the criterion to ODE systems with additive Gaussian and multiplicative lognormal measurement error, it has the potential to be extended to a broad range of problems, including discrete-time systems, PDE systems, and problems with non-Gaussian noise. Assessing identifiability in PDE systems is a challenging problem, which has been addressed recently by [52, 53, 54], but it still remains difficult (if not impossible) to test structural identifiability for arbitrary PDE systems. Therefore, methods that can efficiently assess the practical identifiability of these systems are particularly important.
@@
Critically, the practical identifiability analyses in this study rely on the efficiency of weak-form parameter estimation. We present a method for weak-form parameter estimation of ODE systems with unobserved compartments, utilizing differential elimination to generate weak-form input-output equations and subsequently applying WENDy. This method is efficient, accurate, and robust to noise. Additionally, applied to a blood-tissue diffusion system (Model (7)), a comparison with the results in [1] yields that the WENDy is better than the Hartley and polynomial test functions and strongly competitive with the interpolated polynomials as well as the pure integral approach. Currently, the optimal choice in test functions for weak-form parameter estimation is an open question. [55] proposes a method for constructing a set of test functions that minimizes the error introduced by quadrature.
@@
Regarding method performance, we note that in all the examples presented above (including at all error levels), WENDy is substantially faster than OE methods. In particular, WENDy is faster than the output error method for both the blood-tissue diffusion and SIR models, and is multiple orders of magnitude faster in the SIR case. Ultimately, the benefit of smaller walltimes for estimation of parameters in both the blood-tissue diffusion and SIR models allows for the calculations on thousands of simulated datasets, like those used to generate Figures 4 and 5. Table 1 demonstrates that for a single error ratio level, the WENDy method is able to generate parameter estimates faster than OE methods, with a particularly large advantage (seconds vs minutes over 1,000 simulated datasets) in the SIR model case. For the SIR model, the accuracy of the weak-form method is the same as the OE method for additive noise ratios between $0\%$ and $15\%$, and more accurate at high noise ratios of $e=20\%$. We further observe that the relative error increases linearly with the additive error ratio for this system. Linear error has also been observed for some systems in [56, 27, 57] when performing equation error regression-based optimization. While the median accuracy of the weak-form method is generally significantly higher than the accuracy of the output error method [28], as observed in the SIR example, we did find that it was slightly lower for the blood-tissue diffusion model. However, in the blood-tissue diffusion example, the weak-form method did not fail to converge, while the OE method (which relies on inspired initial parameter estimates) failed in over half of the simulated datasets. More specifically, for Example 3.1, $\sim 60\%$ of the OE optimizations failed to converge, even with reasonable initial parameter guesses (see Figure 7). While in this study, we compare the computational efficiency with respect to a single instance of parameter estimation, these comparisons are important and scalable to more statistically sophisticated modeling frameworks, which often require repeated applications of a parameter estimation method.
@@
One limitation of this study is the reliance on the generation of the weak-form input-output equation. The generation of the full input-output equation of systems of higher dimension can quickly become computationally difficult. Future work will address weak-form parameter practical identifiability for systems where input-output equations cannot be explicitly found. Additionally, information about the initial conditions of the system is lost in the weak-form transformation. There are examples of ODE systems where initial conditions can play an important role in the identifiability of key model parameters [36, 58, 59]. A possible direction to address this limitation is to work with an integral representation of the equation near the end points of the domain to recover information about the initial condition lost in the transformation to the weak form. Future exploration of the impact of initial conditions and methods to incorporate this lost information in our weak-form framework is needed.
Lastly, in this study, we approach practical identifiability through simulated data, and our $(e,q)$-criterion provides an a priori method of studying model identifiability and robustness to noise. We note that our criterion is local, and that following best modeling practices, one may need to test a range of values in the parameter space rather than near a single parameterization of a system. The (e,q)-identifiability method assesses uncertainty in the parameter estimates through repeated simulation by leveraging the fast computational time of the WENDy method. An alternative local method, the FIM method, estimates the uncertainty without simulated data by leveraging information about the local approximation of the optimization problem. While this method saves computation time and expense, the reliability of the estimate generated from the FIM strongly depends both on the nonlinear system being studied as well as on the state of the system.
@@
Additionally, it is also of interest to find methods of assessing the identifiability given a specific experimental dataset. A future direction of this work could be to explore the relationship between experimental data and the stability of identifiability calculations.